OpenStudy (anonymous):

write 5 problems with addition or subtraction of integers. All of the integers used, including the answers, should be between -10 and 10.

OpenStudy (hihi67):

That's not that hard...

OpenStudy (hihi67):

...?

OpenStudy (anonymous):

@dumbcow

OpenStudy (hihi67):

1+2 = 3 3 - 1 = 2 4 - 7 = -3 3 + 9 = 12 4 - 6 = -2

OpenStudy (hihi67):

Do you get it?

OpenStudy (anonymous):

yup I got some like 5 - 2 = 3 and +5 - -2 = +3

OpenStudy (hihi67):

Yeah, that's right too.

OpenStudy (hihi67):

Do I maybe get a medal? lol

OpenStudy (hihi67):

hmmm? @iceicebaby

OpenStudy (hihi67):

Thanks.

OpenStudy (anonymous):

no problem sorry for the wait I was looking for my phone